A Bibliography of Publications in Biometrics: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

25 June 2020 Version 1.00

Title word cross-reference

(θ_{1} + dθ_{2})/θ_{3} [225]. 2 × 2 [181, 182, 127, 178, 250]. 2 × 2 × 2 [89]. 2 × J × K [151]. 2 × K [223]. K 2 × J [298]. P [250]. Pr(X < Y) [201, 317]. Pr(X_{1} < X_{2}) [206]. R [128]. R^{2} [205].

-values [250].

37 [100]. 38 [135].

40 [178]. 41 [171, 186, 172, 160]. 42 [204, 248, 273, 317, 266, 249]. 43 [331]. 44 [281, 330, 342, 303]. 45 [341].

83m [135, 100]. 86d [178]. 87g [266]. 88c [317]. 89d [342].

90k [341].

ABO [124]. ABO-like [124]. abundance [96]. according [50]. accrual [242]. accumulating [338]. across [141]. action [43]. Addendum [173]. additive [77]. additive-hazards [77]. adequacy [234, 331]. adjusting [130]. adjustment [163, 307]. aerial [307]. age [84, 24, 229]. age-specific [24]. agreement [220, 126]. AIDS [291]. allele [45, 28, 55]. alleles [124]. along [310]. Alternative [201, 317]. among [92, 315]. analogue [128]. Analyses [52, 321]. analysing [150]. Analysis [156, 241, 291, 300, 185, 273, 92, 13, 322, 62, 276, 86, 63, 187, 282, 165, 12, 6, 78, 166, 186, 82, 33, 205, 297, 136, 106, 287, 64, 69, 227, 337, 149, 113, 215, 208, 164, 224, 236, 95, 85, 323, 302, 325, 263, 138, 99]. Analyzing [306, 132, 283, 110]. animal [269]. animals [203]. ANOVA [221, 323]. ANOVA-like [323]. antibiotic [50]. appear [264]. applicability [195]. Application [305, 51, 49, 100, 52, 291, 26, 115, 288, 169]. applications [50, 336, 304, 61, 97]. applied [42]. approach [275, 261, 117, 282, 293, 69, 214, 168, 333, 279, 286, 284, 303]. Approximate [62, 267, 177]. Approximately [210, 182]. arbitrarily [26]. arch [85]. Area [39]. ARMA [299]. aspects [63]. assay [320]. assays [200, 156, 20]. Assessing [133, 294, 252]. Assessment [79, 40, 35]. Associated [193, 192, 194]. association [118, 181, 106, 182, 151]. assuming [146]. Asymptotic [244, 128]. Asynchronous [212]. attacks [4]. Attributable [268]. autocorrelated [164]. autoregressive [146].

based [243, 12, 292, 281, 262, 215, 160, 145, 134, 111]. baselines [241]. Bayes [312, 73]. Bayesian [100, 186, 254, 49, 48, 166, 214, 168, 279, 138]. be [259]. been [141]. bending [44]. Bernoulli [167]. Best [25, 195, 335]. between [212, 244, 327, 294, 118, 106, 149]. bias [79, 181, 140]. Biases [233]. binary [277, 170, 172, 139, 285]. binomial [244, 37, 246, 51, 52, 103, 93, 267, 214, 250]. bioassay [117, 288, 304, 176, 266, 58, 280]. bioassays [184, 176, 266]. bioequivalence [35, 107]. biological [108]. Biometrics [135, 100, 204, 248, 171, 273, 186, 281, 172, 178, 160, 317, 266, 330, 331, 249, 341, 342, 303]. biopsies [109]. bivariate [202]. block [185, 273, 20]. blocked [65]. blocking [217]. blocks [6]. blood [14]. Bonferroni [232, 218]. bootstrap [306]. both [179]. boundaries [261, 210]. Bounds [272]. Bradley [83]. Branching [155].

C [99, 211]. calculations [253]. can [259, 264]. cancer [138]. Canonical [86, 337]. capability [292]. capture [190, 235, 308, 8, 119, 42, 311]. carcinogenesis [12]. Carcinogenic [40]. carcinogenicity [51, 269, 300]. case [150, 189, 128, 278, 220, 268, 332]. case-control [150, 278, 220, 268, 332]. catchability [235]. categorical [163, 158, 110, 29, 64, 69, 123, 263]. categories [78, 314]. categorized [201, 317]. causation [334]. causes [264]. cautionary [132]. cell [65]. cell-cycle [65]. censored [101, 75, 291, 165, 199, 26, 177, 313, 111, 173, 153]. censoring [314, 76, 258, 326, 341, 342]. Certainty [147]. chain [11, 37].

Change [309, 49, 100, 7]. Change-in-ratio [309, 7]. change-point [49, 100]. changes [231, 258, 326, 341, 342]. character [161]. characterization [50]. chi [70]. Choice [223]. choosing [315]. Circulant [72]. class [321, 297, 313]. classes [17, 88]. classification [50]. classifications [78]. clinical [261, 338, 241, 166, 186, 239, 238, 34, 281, 262, 80, 242, 224, 325, 279, 315, 102, 81]. closed [245]. cluster [215]. clustered [328]. Clustering [32, 251, 120]. Cochran [23, 29]. coefficient [204, 183, 248, 270, 293, 230]. coefficients [254, 136, 217]. cohort [84, 141]. cohorts [233, 332]. collection [127, 178]. column [223]. Combination [58, 176, 266]. Combining [184, 339]. Comments [194]. common [137]. commonly [181]. community [53]. comparative [12]. comparing [255, 131, 250]. Comparison [35, 67, 204, 183, 248, 27, 83, 111, 258, 326, 341, 342]. comparisons [243]. compartment [318]. compartmental [91, 275]. competing [272]. complex [281, 262]. component [227]. components [226, 146, 143]. composite [34]. Computing [5]. conditional [326, 341]. Confidence [301, 142, 238, 112, 225, 62, 244, 327, 246, 255, 204, 183, 248, 222, 265, 206, 314, 337, 312, 119, 288, 111, 122, 167, 280]. Configural [104]. Confounding [234, 331]. conic [85]. consent [240]. consideration [109]. considerations [38]. consistent [179]. Constant [190, 59]. Constant-parameter [190]. constrained [94]. constraints [288]. construction [44]. Contact [19]. contagion [19]. contingency [223, 57, 33, 287, 250]. continuation [102, 81]. continuous [163, 69]. control [150, 278, 220, 268, 332, 315, 234, 331]. controlling [174]. controls [128]. Cornfield [15]. Correction [135, 100, 103, 204, 248, 171, 273, 55, 186, 30, 232, 281, 172, 178, 160, 317, 266, 330, 102, 331, 249, 341, 342, 303]. Corrections [101, 267]. Correlated [285, 276]. correlation [294, 204, 183, 248, 82, 217]. correlations [144, 257]. cost [182]. cost-effective [182]. count [260, 330]. Counting [170, 39]. covariance [105, 63, 241, 6, 329, 197, 64, 299, 263]. covariate [163, 172, 139]. covariates [296, 136, 112, 285, 73]. Cox [135, 56, 112, 13]. criteria [32]. criterion [251]. Cross [227, 1, 78]. cross-classifications [78]. Cross-validation [227]. crosses [72, 1]. crossing [173, 153]. crossover [130, 166, 186, 237]. cryptic [45]. cultivars [174]. Curve [116, 117, 105, 198, 73, 230]. curves [136, 134]. cycle [65].

D [99, 211]. Data [99, 211, 3, 74, 338, 86, 235, 308, 101, 75, 277, 52, 291, 165, 199, 300, 26, 185, 273, 137, 265, 110, 296, 321, 29, 64, 144, 251, 268, 320, 119, 123, 214, 177, 313, 108, 179, 305, 120, 201, 317, 111, 257, 173, 153, 260, 330, 77, 230, 286, 303]. death [114, 264]. decision [261, 35]. decision-making [35]. decision-theoretic [261]. degree [336]. deletion [162]. density [196, 134]. dependence [78, 172, 139]. dependent [277, 272, 264, 311]. Design [325, 130, 6, 166, 186, 274, 240, 143, 208, 279, 315]. Designs [200, 332, 114, 324, 185, 273, 209, 237, 146, 20]. desirability [130]. Detectability [216]. detecting [259, 231]. detection [49, 100, 310]. determination [327, 239]. determinations [90]. determining [251, 250]. diagnostic [79, 339, 54]. diallel [72, 1]. difference [244, 327]. different [16, 35]. differential [133]. digesta [318]. diploid [28, 55]. discrete [118]. discriminant [187, 69, 302]. Discussion [193]. disease [79, 106]. Diseases [17, 37, 115]. disequilibrium [293]. dispersion [202]. displacement [161]. distance [212, 149, 313]. distributed [239]. Distribution [206, 314, 155, 88, 292, 76, 182, 149, 202, 169, 229, 290, 97]. Distribution-free [206, 314]. Distributional [256]. distributions [103, 93, 31, 96]. diversity [256, 97]. DNA [212]. dose [36, 333, 289, 207, 191, 249]. dose-response [207]. Doses [147, 48]. double [203]. double-sampling [203]. doubling [65]. doubly [291]. doubly-censored [291]. Drug [147, 133]. drugs [43]. due [140]. duration [242].

each [17, 285]. Early [103, 93]. effect [133, 137, 295, 304]. effective [48, 182, 247]. effects [130, 277, 259, 84, 226, 83, 168, 316]. Efficiencies [114]. Efficiency [66, 189, 243, 31, 38, 240]. Efficient [230, 209, 65]. elimination [115]. elliptic [134]. emphasis [91, 239]. Empirical [312, 73]. endpoints [131, 224]. entropy [53]. environment [113]. epidemic [11]. epidemiologic [209, 38, 234, 331]. equal [143]. equation [236, 286, 303]. equilibrium [27, 9]. equivalence [123]. error [302]. errors [276, 146, 164]. establish [123]. estimate [270, 219]. estimated [265, 280]. estimates [198, 44, 289]. Estimating [203, 245, 235, 308, 16, 269, 271, 213, 14, 90, 109, 286, 303, 45]. Estimation [87, 137, 302, 257, 2, 138, 258, 326, 341, 342, 91, 39, 17, 8, 31, 267, 76, 84, 237, 268, 96, 146, 340, 143, 242, 299, 305, 201, 317, 229, 175, 154, 65, 73, 188, 191, 249]. estimator [189, 171, 152, 329, 151]. Estimators [219, 179, 196, 298, 309, 195, 36, 25, 134]. Evaluating [220, 54]. Evaluation [209]. Exact [51, 278, 226, 123, 122, 128, 89, 250, 280]. example [132]. expected [259, 5, 225, 323]. experiment [48, 300, 340, 311]. experimental [114, 315]. experiments [308, 180, 269, 12, 312, 83, 299, 65]. Exploratory [99]. Exploring [211]. exponential [198]. exponentially [239]. exposure [150]. Extended [213]. extension [74, 7]. extensions [158]. extinction [18]. extrapolation [333]. extreme [61].

F [99, 211]. factor [328]. factors [264]. failure [165, 199, 13, 333]. false [109]. familial [144, 257]. families [125]. family [246, 113, 77]. fathers [45]. fertilities [305]. Fieller [266, 176]. finite [18, 67]. fit [56, 135, 328]. fitted [136]. Fitting [108, 198, 116]. fixed [45, 189]. fixed-strata [189]. flow [318]. fluctuation [52]. flushing [134]. follow [137]. follow-up [137]. followed [141]. following [301, 238, 122]. Foss [198]. free [206, 314]. freedom [336]. frequencies [78, 175, 138]. frequency [292, 46, 141]. Friedman [29]. function [112]. functions [255, 149, 336, 216].

G [22]. G. [21]. Gemmell [23]. gene [113, 46, 175, 174]. gene-environment [113]. gene-for-gene [174]. gene-frequency [46]. genealogy [228]. General [43, 126, 37, 157, 222, 297]. generalization [87, 151]. Generalized [3, 298, 36, 124, 252, 318, 306, 336, 253, 286, 303]. generation [155]. generation-time [155]. genes [174]. genetic [124, 30, 10, 44, 121, 305, 61]. genotype [106]. genotypes [28, 55]. geographic [46]. goodness [56, 135, 328]. goodness-of-fit [328]. Group [102, 81, 301, 241, 324, 238, 120, 122, 210, 234, 331]. group-specific [120]. grouped [163, 74]. groups [251, 126, 120]. growth [117, 105, 136, 73, 230]. growth-curve [117, 105, 73]. guide [40].

habitable [118]. habitat [309]. Haenszel [171, 189, 152, 298, 329, 213, 151, 219, 179]. Hardy [27, 293, 9]. Hartley [68]. harvesting [24]. having [17, 78, 263]. hazard [112]. hazards [74, 199, 129, 295, 173, 153, 77]. heritability [142]. Herman [68]. higher [136]. higher-order [136]. Hoaglin [99, 211]. homogeneity [295]. homologous [212]. hospitalization [149]. host [174]. hypotheses [144, 34]. hypothesis [69]. hypothesis-testing [69].

identified [335]. II [279]. III [175]. illness [114]. illness-death [114]. improved [232, 218]. Inbreeding [125]. incidence [269, 109]. including [157]. Incomplete [20, 283, 185, 273, 299, 140]. Incorporating [287, 207]. Increasing [147, 333]. independence [328, 223, 57]. index [30, 10, 202, 256]. indices [44, 97]. indirect [14]. individual [126, 334, 305, 73]. inequalities [61]. inequality [288]. infectious [37]. inference [254, 48, 170, 278, 195, 230, 311]. inferring [174]. influence [252]. information [1, 287]. informative [258, 326, 341, 342]. initial [198]. integration [198]. interaction [89, 113]. interclass [144, 257]. interpretation [205]. interrater [220]. Interval [340, 191, 249, 183, 248, 165, 199, 267, 34, 154, 67, 260, 330, 280]. interval-censored [165, 199]. intervals [62, 244, 327, 246, 301, 204, 222, 265, 206, 314, 142, 238, 312, 112, 225, 111, 122, 167]. intervened [169]. intervention [247]. intraclass [204, 183, 248, 57, 257]. invariant [96]. inverse [245]. Investigating [148, 120]. irregular [116]. ischemic [4]. isolates [174].

J [99, 211, 104]. J. [22]. Jackknife [270]. Jerome [15]. joint [43]. Jolly [87, 119]. judgment [2].

Kaplan [98, 313]. Kendall [217]. Kernel [289]. kinetics [207]. known [316]. Kolmogorov [26].

Large [159, 189, 179]. large-sample [189]. large-strata [179]. Least [159, 13]. Least-Squares [159]. Lefort [21]. length [242, 229]. lengths [160, 145]. lethal [191, 249]. levels [51]. life [56, 135, 62, 255, 101, 75, 71]. life-table [62]. like [124, 323]. likelihood [17, 105, 96, 146, 143, 121, 177, 299, 305, 229, 42, 260, 330, 290, 284]. likelihoods [228]. limiting [179]. limits [119, 288]. Line [310, 156, 134, 20]. line-transect [134]. linear [158, 254, 47, 222, 78, 142, 321, 195, 106, 252, 306, 25, 336, 253, 263, 326, 341]. lines [67]. link [336]. linkage [121]. literature [40]. litter [208]. litter-matched [208]. Locally [259]. locate [30, 10]. locus [60]. Log [106, 78, 329, 281, 262, 219]. log-linear [78]. Log-linear-model [106]. log-odds-ratio [329, 219]. log-rank [281, 262]. Logarithmic [221]. logistic [213]. logit [157]. longitudinal [321, 260, 330, 286, 303]. losing [28, 55]. losses [148]. low [333]. low-dose [333].

made [116]. main [226]. make [264]. making [35, 46]. manifestation [17]. Mann [314]. Mantel [171, 189, 152, 298, 329, 213, 151, 219, 179]. many [120]. maps [46]. marginal [329]. Mark [59, 340]. Mark-recapture [59, 340]. Markov [172, 139, 284]. matched [278, 268, 208]. Matching [38, 270, 140, 141]. matrices [6, 197, 143, 263]. matrix [105]. Maximally [70]. Maximum [96, 146, 143, 299, 229, 42, 17, 148, 305]. McNemar [128]. mean [327, 255, 301, 101, 75, 5, 225]. means [327]. measure [182, 234, 331]. measurements [283, 282, 116, 66, 322]. measures [82, 197, 181, 146, 306, 126, 299]. measuring [149]. median [111, 191, 249]. medical [169]. Meier [98, 313]. Menten [236]. metabolic [115]. method [150, 198, 115, 215, 14, 7, 174, 109]. methodology [71]. methods [132, 283, 180, 183, 248, 170, 64, 35, 305, 307]. Michaelis [236]. minimum [174]. missing [29]. misspecified [129]. mitosis [65]. Mixed [260, 330, 187, 110, 142, 143]. mixed-model [110]. mixed-variable [187]. mixing [31]. mixture [31, 290]. Mixtures [69, 43, 319, 32]. model [56, 135, 163, 74, 11, 37, 158, 114, 87, 254, 117, 105, 157, 277, 47, 4, 165, 199, 222, 110, 142, 106, 226, 83, 320, 112, 13, 119, 143, 295, 24, 250, 334, 42, 174, 207, 326, 341]. modeling [258, 342]. Modelling [296]. Models [286, 303, 91, 43, 275, 190, 78, 321, 197, 195, 59, 129, 252, 213, 318, 306, 172, 139, 336, 168, 179, 304, 253, 260, 330, 263, 77, 230, 284]. Modification [44]. Modified [26, 29]. moments [155]. Monitoring [338]. monotone [101, 75]. Morgenthaler [104]. mortality [149, 138]. mosaics [170]. most [259]. Mosteller [99, 211]. moves [256]. MR0675315 [100]. MR0677963 [135]. MR0786176 [178]. MR0845630 [266]. MR0872965 [317]. MR0931633 [342]. MR1029611 [341]. multilocus [95]. multinomial [88]. multiple [245, 51, 157, 243, 300, 12, 92, 131, 224]. multiple-recapture [245]. multiplicity [214]. Multivariate [6, 176, 266, 163, 117, 156, 105, 76, 213, 184, 143, 257, 32, 58, 167, 280]. mutagenicity [52].

Natural [193, 45, 192, 175, 194]. nearest [41]. negatives [109]. neighbours [41]. Nested [164]. net [272]. networks [19]. no [135, 100, 204, 248, 171, 273, 186, 89, 281, 172, 178, 160, 317, 266, 330, 331, 249, 341, 342, 303]. nonadditivity [336]. nonagreements [92]. nondecreasing [283]. noninteractive [157]. nonlinear [91, 168, 24, 304, 207]. Nonparametric [283, 180, 76, 337, 242, 48, 297, 322]. nonreplicated [265]. nonstratified [148]. normal [301, 31, 182, 34, 32]. normals [290]. note [8, 198, 329, 14]. nucleotide [292]. null [290]. number [45, 88, 251, 256, 14, 174]. numerical [198].

Obituary [21, 15, 22, 68]. observation [285]. observational [209, 316]. Observations [320, 328, 29]. observed [88]. observer [126]. observer-agreement [126]. observing [216]. Obtaining [119, 198]. occupancy [94]. occur [231]. occurrence [4, 292]. odds [3, 171, 152, 78, 33, 329, 219]. offspring [45]. onchocerciasis [109]. One [261, 336, 210]. one-parameter [210]. One-sided [261]. only [259]. open [332]. optimal [274, 210]. Optimum [24]. order [132, 136, 61]. ordered [163, 78, 223, 314, 110, 29, 123, 2]. ordinal [3, 150]. organisms [16]. Otto [68]. outlier [47]. overlap [292]. ovulation [49, 100].

paired [296, 83]. paired-comparison [83]. pairs [118]. parallel [156, 241, 20]. parallel-group [241]. parallel-line [156, 20]. Parameter [91, 190, 137, 314, 210]. parameters [244, 267, 265, 44, 59, 96, 119, 288, 73]. Parental [228]. partial [72, 252, 151]. particles [14]. Path [113]. pathogen [174]. pathogens [50]. patients [259]. patterned [105]. patterns [132, 50]. percentage [290]. performance [121]. period [166, 186, 84]. personal [23]. Phase [325, 279]. plankton [16]. Planning [65, 182]. plans [72, 175]. plant [174]. plots [215]. point [132, 39, 49, 100, 196, 154, 216]. point-counting [39]. points [116, 290]. Poisson [330, 202, 169, 260]. polynomial [136]. Polysampling [104]. pooling [141]. population [203, 327, 235, 308, 16, 18, 309, 231, 271, 340, 134, 24, 305, 65]. population-doubling [65]. populations [45, 245, 175, 307]. possible [124]. Post [80]. Post-stratification [80]. potency [280]. Power [127, 178, 253, 128, 148]. Power/sample [253]. powerful [259]. Practical [104]. prediction [49, 100]. predictive [339]. presence [124, 76, 217, 258, 326, 341, 342]. pretreatment [241]. prevalent [233]. price [98]. principal [227]. prior [287]. probabilities [62, 272, 311]. Probability [18, 45, 28, 55, 250, 334, 310]. probit [157]. Probits [319]. Problem [159, 49, 100, 274, 149, 214, 61]. problems [94]. procedure [115, 110, 232, 218, 213, 177, 311]. Procedures [131, 26, 201, 317]. process [155, 258, 342]. processes [196, 294, 216]. progressive [314]. projections [215]. Properties [129, 30, 10, 256, 97]. proportion [203, 31, 231, 90]. proportional [74, 199, 129, 112, 295, 77]. proportional- [77]. proportional-hazard [112]. proportional-hazards [74, 129]. proportions [246, 17, 250]. protective [264]. purity [90]. quantal [117, 48, 320, 288]. quantal-response [48]. quantile [215]. quantile-quantile [215]. quantiles [62]. quasi [284]. quasi-likelihood [284].

Random [83, 304, 277, 226, 168, 230]. random-coefficient [230]. Random-effect [304]. Randomization [240, 148, 322]. randomized [6, 80]. randomly [101, 75]. rank [231, 281, 262]. ranked [2]. ranks [108]. rate [333]. raters [92]. Rates [193, 192, 87, 269, 84, 329, 302, 138, 194, 109]. ratio [246, 200, 105, 171, 152, 267, 142, 309, 329, 225, 121, 7, 219, 290]. ratios [3, 254, 33]. reactions [174]. recapture [245, 190, 235, 308, 8, 59, 119, 340, 42, 311]. recessive [124]. recombination [95]. reconstruction [228]. reduction [328]. reflected [111]. Reflexive [41]. regions [337, 102, 81]. regression [56, 135, 165, 205, 129, 213, 295, 285, 304, 67, 260, 330, 284]. rejective [232, 218]. related [125, 312]. relationship [73]. relative [189, 309, 36, 280]. removal [154]. reparameterization [30, 10]. Repeated [208, 283, 282, 82, 116, 197, 66, 146, 306, 322, 299]. repeated-measures [82, 197]. resemblance [113]. residual [130, 255, 101, 75]. resighting [87]. respond [259]. Response [147, 48, 289, 207]. responses [156, 239]. restricted [103, 93]. results [155, 31, 58]. Review [99, 211, 104, 267, 64]. rhythm [108]. rhythmometry [274]. right [26, 258, 326, 341, 342]. right-censored [26]. risk [40, 268, 334, 264]. risks [272]. Robust [99, 196]. robustness [66, 104]. rodent [300]. role [71, 316]. Route [104]. rules [339]. run [160, 145]. runs [160, 145].

S [104]. sacrifice [180]. Sample [327, 239, 89, 281, 262, 16, 189, 53, 329, 231, 182, 271, 14, 288, 127, 178, 253, 173, 153, 97, 167]. sampled [28, 55]. samples [244, 87, 277, 131, 2]. sampling [203, 245, 30, 10, 220, 175, 310]. sciences [71]. Score [295, 124, 129]. scores [223]. screening [115]. screens [51]. Seber [87, 119]. second [132]. second-order [132]. sections [85]. seed [90]. seed-purity [90]. selected [70]. Selecting [1]. Selection [187, 79, 125, 18, 30, 10, 44, 302, 61, 60]. semiparametric [165]. sensitive [173, 153]. sequences [212, 172, 139]. Sequential [34, 261, 301, 103, 93, 324, 115, 238, 122, 210, 102, 81, 175]. sequentially [232, 218]. serial [82]. series [312, 284]. set [251, 2]. sets [255, 86]. several [86, 315]. severity [17]. shape [85]. Shapes [211]. shared [323]. sib [90, 228]. sided [261]. Sightability [307]. significance [51, 294, 123, 214, 208]. similarity [270, 323]. simple [339]. Simplified [154]. Simulated [290]. simulation [243]. simulation-based [243]. Simultaneous [222, 167]. single [47, 271, 240, 60]. single-consent [240]. singly [173, 153]. singular [263]. sinusoid [108]. size [327, 245, 235, 308, 239, 309, 271, 340, 127, 178, 253]. sizes [16, 89, 281, 262]. Skellam [22]. skewness [267]. skin [109]. skin-snip [109]. slope [200]. Small [288, 244, 231, 182, 340, 14, 120, 138, 167]. Small-sample [288, 182, 167]. Smirnov [26]. smoothers [322]. snip [109]. Solution [159]. Some [158, 196, 63, 94, 31, 318, 61, 97, 64]. Sparse [159, 308, 137, 179]. spatial [196, 294, 265, 195, 120, 216]. Spearman [36, 217]. species [118, 53, 323, 174, 97]. species-diversity [97]. specific [87, 285, 24, 120]. spectral [321]. splines [188]. square [70]. Squares [159, 5, 251, 225, 13]. stable [182]. stage [315]. statistic [281, 262]. Statistical [195, 236, 95, 85, 147, 94, 50, 53, 64]. Statistics [45, 193, 175, 192, 71, 70, 313, 61, 194]. stereological [195]. stereology [195, 25]. stochastic [4, 320, 318, 334]. stopping [261, 103, 93]. strata [189, 179, 141]. stratification [80]. stratified [33]. structure [299]. Structured [159, 197]. structures [195]. studies [150, 278, 209, 182, 38, 148, 332, 316, 234, 331]. study [19, 321, 242, 141]. subject [79]. subjects [126]. subsequences [292]. subset [259]. sum [251]. sum-of-squares [251]. surveys [307]. Survival [188, 62, 87, 291, 180, 296, 297, 272, 112, 208, 148, 313, 111, 77]. survival/sacrifice [180]. susceptibility [50]. Symmetric [324]. synthetic [332]. systems [124].

table [62]. Tables [211, 223, 298, 57, 89, 33, 181, 287, 151, 127, 178, 250]. tag [87]. tag-resighting [87]. techniques [39]. terms [155]. Terry [83]. test [301, 105, 52, 128, 324, 26, 124, 50, 89, 82, 232, 218, 29, 202, 121, 160, 145, 111, 173, 153, 290, 122]. test-based [111]. Testing [56, 135, 276, 47, 57, 144, 335, 9, 247, 147, 60, 51, 223, 293, 69, 123, 214, 208, 46, 107]. Tests [255, 101, 75, 161, 162, 79, 328, 259, 103, 93, 94, 27, 231, 297, 226, 238, 129, 339, 295, 313, 322, 54]. their [292, 73, 97]. theorem [176, 266]. theoretic [261]. time [87, 155, 165, 199, 116, 59, 13, 111, 65, 311, 284]. time-dependent [311]. time-points [116]. time-specific [87]. times [12]. tolerance [337]. total [242]. toxicity [157]. transect [134, 310]. transformations [221, 157]. transient [4]. trapping [154]. treated [231]. treatment [259, 335, 123]. treatments [315]. Trends [211]. trial [338, 80]. trials [261, 241, 166, 186, 239, 238, 34, 281, 262, 242, 224, 325, 279, 315, 210, 102, 81]. triangular [102, 81]. tribute [23]. truncation [18]. Tukey [99, 211, 104]. tumor [269, 300, 12]. Two [231, 244, 327, 246, 255, 17, 294, 31, 166, 186, 226, 339, 67, 173, 153, 315, 290]. two-period [166, 186]. Two-sample [231, 173, 153]. two-stage [315]. two-way [226]. type [314]. types [16].

Unbalanced [197, 142, 226, 230]. unbiased [195, 25]. uncatchable [203]. uncensored [173, 153]. Understanding [99]. unequal [235, 6]. unified [275]. units [118]. univariate [82, 120]. unknown [310]. usage [309]. use [132, 244, 205, 66, 136]. used [181]. using [245, 101, 75, 1, 78, 251, 34, 83, 112, 13, 208, 108, 322, 323, 2, 188]. validation [227]. validity [38]. value [339]. values [250]. Variability [193, 275, 192, 194]. variable [187, 217]. variables [163, 158, 150, 187, 69, 302, 263]. variance [276, 171, 152, 6, 142, 226, 182, 237, 146, 143, 164, 219, 179, 2]. variate [86, 337]. variates [34]. verification [79]. versus [149, 111]. via [306]. Vital [193, 192, 84, 194]. Vital-Rates [193].

W [99, 211, 104]. way [226]. Weibull [157]. Weighted [313, 328, 13]. Weinberg [27, 293, 9]. whether [335]. Whitney [314]. Wilcoxon [314]. wildlife [307]. William [23]. women [49, 100].

Zelen [240].

References

Curnow:1980:SCU

[1] R. N. Curnow. Selecting crosses using information from a diallel cross. Biometrics, 36(1):1–8, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Stokes:1980:EVU

[2] S. L. Stokes. Estimation of variance using judgment ordered ranked set samples. Biometrics, 36(1):35–42, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Agresti:1980:GOR

[3] Alan Agresti. Generalized odds ratios for ordinal data. Biometrics, 36 (1):59–67, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dunn:1980:SMO

[4] Judith K. Dunn and Robert J. Hardy. A stochastic model for the occurrence of transient ischemic attacks. Biometrics, 36(1):91–103, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goodnight:1980:CEM

[5] J. H. Goodnight and F. M. Speed. Computing expected mean squares. Biometrics, 36(1):123–125, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

GarciaBen:1980:MAV

[6] M. S. Garcia Ben and V. J. Yohai. Multivariate analysis of variance in a randomized blocks design when covariance matrices are unequal. Biometrics, 36(1):127–133, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Otis:1980:ECR

[7] David L. Otis. An extension of the change-in-ratio method. Biometrics, 36 (1):141–147, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Darroch:1980:NCR

[8] J. N. Darroch and D. Ratcliff. A note on capture–recapture estimation. Biometrics, 36(1):149–153, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ledwina:1980:THW

[9] Teresa Ledwina and Stanisław Gnot. Testing for Hardy–Weinberg equilibrium. Biometrics, 36(1):161–165, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hayes:1980:RGS

[10] J. F. Hayes and W. G. Hill. A reparameterization of a genetic selection index to locate its sampling properties. Biometrics, 36(2):237–248, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [30].

Becker:1980:ECM

[11] Niels Becker. An epidemic chain model. Biometrics, 36(2):249–254, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gail:1980:ACC

[12] M. H. Gail, T. J. Santner, and C. C. Brown. An analysis of comparative carcinogenesis experiments based on multiple times to tumor. Biometrics, 36(2):255–266, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lubin:1980:AUC

[13] Jay H. Lubin. Analysis under Cox’s failure time model using weighted least squares. Biometrics, 36(2):307–312, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Minder:1980:NIM

[14] Ch. E. Minder. A note on an indirect method for estimating the number of small particles in a blood sample. Biometrics, 36(2):313–316, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1980:OJC

[15] Anonymous. Obituary: Jerome Cornfield, 1912–1979. Biometrics, 36(2): 357–358, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dahiya:1980:EPS

[16] Ram C. Dahiya. Estimating the population sizes of different types of organisms in a plankton sample. Biometrics, 36(3):437–446, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonaiti:1980:DHT

[17] Catherine Bona¨ ıti and Florence Demenais. Diseases having two classes of severity: estimation of the proportions of each manifestation by maximum likelihood. Biometrics, 36(3):447–460, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ghai:1980:PEF

[18] G. L. Ghai. Probability of extinction in a finite population with truncation selection. Biometrics, 36(3):461–472, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hartigan:1980:CNS

[19] P. M. Hartigan. Contact networks and the study of contagion. Biometrics, 36(3):473–485, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Win:1980:IBD

[20] Kyi Win and A. Dey. Incomplete block designs for parallel-line assays. Biometrics, 36(3):487–492, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1980:OGL

[21] Anonymous. Obituary: G. Lefort, 1921–1979. Biometrics, 36(3):557–558, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1980:OJG

[22] Anonymous. Obituary: J. G. Skellam, 1914–1979. Biometrics, 36(3): 558–559, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anderson:1980:WGC

[23] R. L. Anderson. William Gemmell Cochran, 1909–1980: a personal tribute. Biometrics, 36(4):574–578, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Reed:1980:OAS

[24] William J. Reed. Optimum age-specific harvesting in a nonlinear population model. Biometrics, 36(4):579–593, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Orive:1980:BLU

[25] Luis-M. Cruz Orive. Best linear unbiased estimators for stereology. Biometrics, 36(4):595–605, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fleming:1980:MKS

[26] Thomas R. Fleming, Judith R. O’Fallon, Peter C. O’Brien, and David P. Harrington. Modified Kolmogorov–Smirnov test procedures with application to arbitrarily right-censored data. Biometrics, 36(4):607–625, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Emigh:1980:CTH

[27] Ted H. Emigh. A comparison of tests for Hardy–Weinberg equilibrium. Biometrics, 36(4):627–642, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gregorius:1980:PLA

[28] Hans-Rolf Gregorius. The probability of losing an allele when diploid genotypes are sampled. Biometrics, 36(4):643–652, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [55].

Klotz:1980:MCF

[29] Jerome Klotz. A modified Cochran–Friedman test with missing observations and ordered categorical data. Biometrics, 36(4):665–670, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hayes:1980:CRG

[30] J. F. Hayes and W. G. Hill. Correction to: “A reparameterization of a genetic selection index to locate its sampling properties”. Biometrics, 36 (4):752, 1980. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [10].

Ganesalingam:1981:SER

[31] S. Ganesalingam and G. J. McLachlan. Some efficiency results for the estimation of the mixing proportion in a mixture of two normal distributions. Biometrics, 37(1):23–33, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Symons:1981:CCM

[32] M. J. Symons. Clustering criteria and multivariate normal mixtures. Biometrics, 37(1):35–43, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heilbron:1981:ARO

[33] David C. Heilbron. The analysis of ratios of odds ratios in stratified contingency tables. Biometrics, 37(1):55–66, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lachin:1981:SCT

[34] John M. Lachin. Sequential clinical trials for normal variates using interval composite hypotheses. Biometrics, 37(1):87–101, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mandallaz:1981:CDM

[35] Daniel Mandallaz and Jochen Mau. Comparison of different methods for decision-making in bioequivalence assessment. Biometrics, 37(2):213–222, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morton:1981:GSE

[36] R. Morton. Generalized Spearman estimators of relative dose. Biometrics, 37(2):223–233, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Becker:1981:GCB

[37] Niels Becker. A general chain binomial model for infectious diseases. Biometrics, 37(2):251–258, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kupper:1981:MES

[38] Lawrence L. Kupper, John M. Karon, David G. Kleinbaum, Hal Morgenstern, and Donald K. Lewis. Matching in epidemiologic studies: validity and efficiency considerations. Biometrics, 37(2):271–291, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bellhouse:1981:AEP

[39] D. R. Bellhouse. Area estimation by point-counting techniques. Biometrics, 37(2):303–312, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krewski:1981:CRA

[40] Daniel Krewski and Charles Brown. Carcinogenic risk assessment: a guide to the literature. Biometrics, 37(2):353–366, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cox:1981:RNN

[41] Trevor F. Cox. Reflexive nearest neighbours. Biometrics, 37(2):367–369, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sprott:1981:MLA

[42] D. A. Sprott. Maximum likelihood applied to a capture–recapture model. Biometrics, 37(2):371–375, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ashford:1981:GMJ

[43] J. R. Ashford. General models for the joint action of mixtures of drugs. Biometrics, 37(3):457–474, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hayes:1981:MEP

[44] J. F. Hayes and W. G. Hill. Modification of estimates of parameters in the construction of genetic selection indices (“bending”). Biometrics, 37 (3):483–493, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Arnold:1981:SNP

[45] Jonathan Arnold. Statistics of natural populations. I. Estimating an allele probability in cryptic fathers with a fixed number of offspring. Biometrics, 37(3):495–504, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Piazza:1981:MTG

[46] A. Piazza, P. Menozzi, and L. Cavalli-Sforza. The making and testing of geographic gene-frequency maps. Biometrics, 37(4):635–659, 1981. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Doornbos:1981:TSO

[47] R. Doornbos. Testing for a single outlier in a linear model. Biometrics, 37 (4):705–711, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Disch:1981:BNI

[48] Damon Disch. Bayesian nonparametric inference for effective doses in a quantal-response experiment. Biometrics, 37(4):713–722, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carter:1981:BCP

[49] Randy L. Carter and Barry J. N. Blight. A Bayesian change-point problem with an application to the prediction and detection of ovulation in women. Biometrics, 37(4):743–751, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [100].

Grimson:1981:STC

[50] Roger C. Grimson and Katherine H. Chavigny. A statistical test for classification, with applications to the characterization of pathogens according to antibiotic susceptibility patterns. Biometrics, 37(4):753–761, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brown:1981:ESL

[51] Charles C. Brown and Thomas R. Fears. Exact significance levels for multiple binomial testing with application to carcinogenicity screens. Biometrics, 37(4):763–774, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Collings:1981:ABD

[52] Bruce J. Collings, Barry H. Margolin, and Gary W. Oehlert. Analyses for binomial data, with application to the fluctuation test for mutagenicity. Biometrics, 37(4):775–794, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holgate:1981:SES

[53] P. Holgate. The statistical entropy of a sample from a community of species. Biometrics, 37(4):795–799, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thibodeau:1981:EDT

[54] L. A. Thibodeau. Evaluating diagnostic tests. Biometrics, 37(4):801–804, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gregorius:1981:CPL

[55] H.-R. Gregorius. Correction to: “The probability of losing an allele when diploid genotypes are sampled”. Biometrics, 37(4):875, 1981. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [28].

Andersen:1982:TGF

[56] Per Kragh Andersen. Testing goodness of fit of Cox’s regression and life model. Biometrics, 38(1):67–77, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [135].

Haber:1982:TII

[57] Michael Haber. Testing for independence in intraclass contingency tables. Biometrics, 38(1):93–103, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Volund:1982:CMB

[58] Aage Vølund. Combination of multivariate bioassay results. Biometrics, 38(1):181–190, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jolly:1982:MRM

[59] G. M. Jolly. Mark-recapture models with parameters constant in time. Biometrics, 38(2):301–321, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Watterson:1982:TSS

[60] G. A. Watterson. Testing selection at a single locus. Biometrics, 38(2): 323–331, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tong:1982:SAI

[61] Y. L. Tong. Some applications of inequalities for extreme order statistics to a genetic selection problem. Biometrics, 38(2):333–339, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anderson:1982:ACI

[62] J. R. Anderson, L. Bernstein, and M. C. Pike. Approximate confidence intervals for probabilities of survival and quantiles in life-table analysis. Biometrics, 38(2):407–416, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cox:1982:SAA

[63] D. R. Cox and P. McCullagh. Some aspects of analysis of covariance. Biometrics, 38(3):541–561, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Koch:1982:RSS

[64] Gary G. Koch, Ingrid A. Amara, Gordon W. Davis, and Dennis B. Gillings. A review of some statistical methods for covariance analysis of categorical data. Biometrics, 38(3):563–595, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staudte:1982:PBM

[65] Robert G. Staudte. Planning blocked mitosis experiments for efficient estimation of population-doubling time and cell-cycle time. Biometrics, 38(3):777–785, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jensen:1982:ERU

[66] D. R. Jensen. Efficiency and robustness in the use of repeated measurements. Biometrics, 38(3):813–825, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Spurrier:1982:CTR

[67] John D. Spurrier, John E. Hewett, and Zuhdi Lababidi. Comparison of two regression lines over a finite interval. Biometrics, 38(3):827–836, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anonymous:1982:OHO

[68] Anonymous. Obituary: Herman Otto Hartley, 1912–1980. Biometrics, 38 (3):855–856, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krzanowski:1982:MCC

[69] W. J. Krzanowski. Mixtures of continuous and categorical variables in discriminant analysis: a hypothesis-testing approach. Biometrics, 38(4): 991–1002, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miller:1982:MSC

[70] Rupert Miller and David Siegmund. Maximally selected chi square statistics. Biometrics, 38(4):1011–1016, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dolby:1982:RSM

[71] G. R. Dolby. The role of statistics in the methodology of the life sciences. Biometrics, 38(4):1069–1083, 1982. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Arya:1983:CPP

[72] A. S. Arya. Circulant plans for partial diallel crosses. Biometrics, 39 (1):43–52, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Strenio:1983:EBE

[73] Judith F. Strenio, Herbert I. Weisberg, and Anthony S. Bryk. Empirical Bayes estimation of individual growth-curve parameters and their relationship to covariates. Biometrics, 39(1):71–86, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aranda-Ordaz:1983:EPH

[74] Francisco J. Aranda-Ordaz. An extension of the proportional-hazards model for grouped data. Biometrics, 39(1):109–117, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chen:1983:TMM

[75] Yuan-Yan Chen, Myles Hollander, and Naftali A. Langberg. Tests for monotone mean residual life, using randomly censored data. Biometrics, 39(1):119–127, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [101].

Hanley:1983:NEM

[76] James A. Hanley and Milton N. Parnes. Nonparametric estimation of a multivariate distribution in the presence of censoring. Biometrics, 39(1):

129–139, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tibshirani:1983:FPA

[77] Robert J. Tibshirani and Antonio Ciampi. A family of proportionaland additive-hazards models for survival data. Biometrics, 39(1):141–147, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Goodman:1983:ADC

[78] Leo A. Goodman. The analysis of dependence in cross-classifications having ordered categories, using log-linear models for frequencies and log-linear models for odds. Biometrics, 39(1):149–160, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Begg:1983:ADT

[79] Colin B. Begg and Robert A. Greenes. Assessment of diagnostic tests when disease verification is subject to selection bias. Biometrics, 39(1): 207–215, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McHugh:1983:PSR

[80] Richard McHugh and John Matts. Post-stratification in the randomized clinical trial. Biometrics, 39(1):217–225, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Whitehead:1983:GSC

[81] John Whitehead and Irene Stratton. Group sequential clinical trials with triangular continuation regions. Biometrics, 39(1):227–236, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [102].

Hearne:1983:TSC

[82] Erwin M. Hearne III, Gary M. Clark, and John P. Hatch. A test for serial correlation in univariate repeated-measures analysis. Biometrics, 39(1): 237–243, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lancaster:1983:REP

[83] J. F. Lancaster and Dana Quade. Random effects in paired-comparison experiments using the Bradley–Terry model. Biometrics, 39(1):245–249, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holford:1983:EAP

[84] Theodore R. Holford. The estimation of age, period and cohort effects for vital rates. Biometrics, 39(2):311–324, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sampson:1983:SAA

[85] Paul D. Sampson. Statistical analysis of arch shape with conic sections. Biometrics, 39(2):411–423, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Campbell:1983:CVA

[86] N. A. Campbell and J. A. Tomenson. Canonical variate analysis for several sets of data. Biometrics, 39(2):425–435, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brownie:1983:ETS

[87] C. Brownie and D. S. Robson. Estimation of time-specific survival rates from tag-resighting samples: a generalization of the Jolly–Seber model. Biometrics, 39(2):437–453, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Emigh:1983:NOC

[88] Ted H. Emigh. On the number of observed classes from a multinomial distribution. Biometrics, 39(2):485–491, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Haber:1983:SSE

[89] Michael Haber. Sample sizes for the exact test of “no interaction” in 2 × 2 × 2 tables. Biometrics, 39(2):493–498, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Samaniego:1983:ESP

[90] Francisco J. Samaniego and Pere Arús. On estimating the sib proportion in seed-purity determinations. Biometrics, 39(3):563–572, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Allen:1983:PEN

[91] D. M. Allen. Parameter estimation for nonlinear models with emphasis on compartmental models. Biometrics, 39:629–637, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

James:1983:ANA

[92] Ian R. James. Analysis of nonagreements among multiple raters. Biometrics, 39(3):651–657, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dambrosia:1983:ESS

[93] James M. Dambrosia and Samuel W. Greenhouse. Early stopping for sequential restricted tests of binomial distributions. Biometrics, 39(3): 695–710, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [103].

Davies:1983:SST

[94] Paul Davies. Some statistical tests for constrained occupancy problems. Biometrics, 39(3):719–725, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Risch:1983:SAM

[95] N. Risch and K. Lange. Statistical analysis of multilocus recombination. Biometrics, 39(4):949–963, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lo:1983:MLE

[96] Hing P. Lo and Jagannath K. Wani. Maximum likelihood estimation of the parameters of the invariant abundance distributions. Biometrics, 39 (4):977–986, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tong:1983:SDP

[97] Y. L. Tong. Some distribution properties of the sample species-diversity indices and their applications. Biometrics, 39(4):999–1008, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Miller:1983:WPK

[98] Rupert G. Miller, Jr. What price Kaplan–Meier? Biometrics, 39(4): 1077–1081, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Healy:1983:RUR

[99] M. J. R. Healy. Review: Understanding Robust and Exploratory Data Analysis, by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 39(4):1126–1127, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). URL http://www.jstor.org/stable/

2531362.

Carter:1983:CBC

[100] R. L. Carter and B. J. N. Blight. Correction to: “A Bayesian change-point problem with an application to the prediction and detection of ovulation in women” [Biometrics 37 (1981), no. 4, 743–751; MR0675315 (83m:62171)]. Biometrics, 39(4):1137, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [49].

Chen:1983:CTM

[101] Yuan-Yan Chen, Myles Hollander, and Naftali A. Langberg. Corrections of: “Tests for monotone mean residual life, using randomly censored data”. Biometrics, 39(4):1137, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [75].

Whitehead:1983:CGS

[102] John Whitehead and Irene Stratton. Correction to: “Group sequential clinical trials with triangular continuation regions”. Biometrics, 39(4): 1137, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [81].

Dambrosia:1983:CES

[103] J. M. Dambrosia and S. W. Greenhouse. Correction to: “Early stopping for sequential restricted tests of binomial distributions”. Biometrics, 39 (4):1138, 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [93].

Rosenberger:1983:RCP

[104] J. L. Rosenberger. Review: Configural Polysampling: A Route to Practical Robustness, by S. Morgenthaler and J. W. Tukey. Biometrics, 39(4):1284, December 1983. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420

(electronic). URL http://www.jstor.org/stable/2532280.

Chinchilli:1984:LRT

[105] Vernon M. Chinchilli and Walter H. Carter, Jr. A likelihood ratio test for a patterned covariance matrix in a multivariate growth-curve model. Biometrics, 40(1):151–156, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Khamis:1984:LLM

[106] H. J. Khamis and K. Hinkelmann. Log-linear-model analysis of the association between disease and genotype. Biometrics, 40(1):177–188, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rocke:1984:TB

[107] David M. Rocke. On testing for bioequivalence. Biometrics, 40(1):225– 230, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pettitt:1984:FSB

[108] A. N. Pettitt. Fitting a sinusoid to biological rhythm data using ranks. Biometrics, 40(2):295–300, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yanagawa:1984:MEI

[109] Takashi Yanagawa, Fumiyoshi Kasagi, and Takesumi Yoshimura. A method for estimating incidence rates of onchocerciasis from skin-snip biopsies with consideration of false negatives. Biometrics, 40(2):301–311, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harville:1984:MMP

[110] David A. Harville and Robert W. Mee. A mixed-model procedure for analyzing ordered categorical data. Biometrics, 40(2):393–408, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Slud:1984:CRV

[111] Eric V. Slud, David P. Byar, and Sylvan B. Green. A comparison of reflected versus test-based confidence intervals for the median survival time, based on censored data. Biometrics, 40(3):587–600, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Link:1984:CIS

[112] Carol L. Link. Confidence intervals for the survival function using Cox’s proportional-hazard model with covariates. Biometrics, 40(3):601–609, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lathrop:1984:PAF

[113] G. Mark Lathrop, J. M. Lalouel, and A. Jacquard. Path analysis of family resemblance and gene-environment interaction. Biometrics, 40(3):611– 625, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Borgan:1984:EED

[114] Ørnulf Borgan, Knut Liestøl, and Peter Ebbesen. Efficiencies of experimental designs for an illness-death model. Biometrics, 40(3):627–638, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Geertsema:1984:ASP

[115] J. C. Geertsema and C. J. Reinecke. The application of a sequential procedure with elimination as a method for the screening for metabolic diseases. Biometrics, 40(3):663–673, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hui:1984:CFR

[116] Siu L. Hui. Curve fitting for repeated measurements made at irregular time-points. Biometrics, 40(3):691–697, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carter:1984:GCM

[117] E. M. Carter and J. J. Hubert. A growth-curve model approach to multivariate quantal bioassay. Biometrics, 40(3):699–706, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Engen:1984:ABP

[118] Steinar Engen. The association between pairs of species on discrete habitable units. Biometrics, 40(3):729–738, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Manly:1984:OCL

[119] Bryan F. J. Manly. Obtaining confidence limits on parameters of the Jolly– Seber model for capture–recapture data. Biometrics, 40(3):749–758, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schweder:1984:IGS

[120] Tore Schweder. Investigating group-specific clustering in univariate spatial data with many small groups. Biometrics, 40(3):767–776, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Nordheim:1984:PLR

[121] Erik V. Nordheim, David M. O’Malley, and Shein-Chung Chow. On the performance of a likelihood ratio test for genetic linkage. Biometrics, 40 (3):785–790, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsiatis:1984:ECI

[122] Anastasios A. Tsiatis, Gary L. Rosner, and Cyrus R. Mehta. Exact confidence intervals following a group sequential test. Biometrics, 40(3): 797–803, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mehta:1984:EST

[123] Cyrus R. Mehta, Nitin R. Patel, and Anastasios A. Tsiatis. Exact significance testing to establish treatment equivalence with ordered categorical data. Biometrics, 40(3):819–825, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gart:1984:STP

[124] John J. Gart and Jun mo Nam. A score test for the possible presence of recessive alleles in generalized ABO-like genetic systems. Biometrics, 40 (4):887–894, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Burrows:1984:IUS

[125] Peter M. Burrows. Inbreeding under selection from related families. Biometrics, 40(4):895–906, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OConnell:1984:GOA

[126] Dianne L. O’Connell and Annette J. Dobson. General observer-agreement measures on individual subjects and groups of subjects. Biometrics, 40 (4):973–983, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Munoz:1984:PSS

[127] Alvaro Muñoz and Bernard Rosner. Power and sample size for a collection of 2×2 tables. Biometrics, 40(4):995–1004, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [178].

Duffy:1984:AEP

[128] Stephen W. Duffy. Asymptotic and exact power for the McNemar test and its analogue with R controls per case. Biometrics, 40(4):1005–1015, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lagakos:1984:PPH

[129] S. W. Lagakos and D. A. Schoenfeld. Properties of proportional-hazards score tests under misspecified regression models. Biometrics, 40(4):1037– 1048, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Abeyasekera:1984:DAR

[130] Savitri Abeyasekera and R. N. Curnow. The desirability of adjusting for residual effects in a crossover design. Biometrics, 40(4):1071–1078, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OBrien:1984:PCS

[131] Peter C. O’Brien. Procedures for comparing samples with multiple endpoints. Biometrics, 40(4):1079–1087, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Baddeley:1984:CEU

[132] A. J. Baddeley and B. W. Silverman. A cautionary example on the use of second-order methods for analyzing point patterns. Biometrics, 40(4): 1089–1093, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berry:1984:ADD

[133] D. A. Berry, M. L. Eaton, B. P. Ekholm, and T. L. Fox. Assessing differential drug effect. Biometrics, 40(4):1109–1115, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Otten:1984:LTE

[134] A. Otten and P. G. de Vries. On line-transect estimators for population density, based on elliptic flushing curves. Biometrics, 40(4):1145–1150, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Andersen:1984:CTG

[135] Per Kragh Andersen. Correction to: “Testing goodness of fit of Cox’s regression and life model” [Biometrics 38 (1982), no. 1, 67–77; MR0677963 (83m:62169)]. Biometrics, 40(4):1217, 1984. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [56].

Kenward:1985:UFH

[136] M. G. Kenward. The use of fitted higher-order polynomial coefficients as covariates in the analysis of growth curves. Biometrics, 41(1):19–28, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenland:1985:ECE

[137] Sander Greenland and James M. Robins. Estimation of a common effect parameter from sparse follow-up data. Biometrics, 41(1):55–68, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tsutakawa:1985:ECM

[138] Robert K. Tsutakawa. Estimation of cancer mortality rates: a Bayesian analysis of small frequencies. Biometrics, 41(1):69–79, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Muenz:1985:MMC

[139] Larry R. Muenz and Lawrence V. Rubinstein. Markov models for covariate dependence of binary sequences. Biometrics, 41(1):91–101, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [172].

Rosenbaum:1985:BDI

[140] Paul R. Rosenbaum and Donald B. Rubin. The bias due to incomplete matching. Biometrics, 41(1):103–116, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Weinberg:1985:PAS

[141] Clarice R. Weinberg. On pooling across strata when frequency matching has been followed in a cohort study. Biometrics, 41(1):117–127, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Harville:1985:CIV

[142] David A. Harville and Alan P. Fenech. Confidence intervals for variance ratio, or for heritability, in an unbalanced mixed linear model. Biometrics, 41(1):137–152, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meyer:1985:MLE

[143] K. Meyer. Maximum likelihood estimation of variance components for a multivariate mixed model with equal design matrices. Biometrics, 41(1): 153–165, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Konishi:1985:THA

[144] Sadanori Konishi. Testing hypotheses about interclass correlations from familial data. Biometrics, 41(1):167–176, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OBrien:1985:RTB

[145] Peter C. O’Brien and Peter J. Dyck. A runs test based on run lengths. Biometrics, 41(1):237–244, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [160].

Mansour:1985:MLE

[146] Hussein Mansour, Erik V. Nordheim, and J. J. Rutledge. Maximum likelihood estimation of variance components in repeated measures designs assuming autoregressive errors. Biometrics, 41(1):287–294, 1985. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1985:TSC

[147] J. W. Tukey, J. L. Ciminera, and J. F. Heyse. Testing the statistical certainty of a response to increasing doses of a drug. Biometrics, 41(1): 295–301, March 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-

0420 (electronic). URL http://www.jstor.org/stable/2530666.

Palta:1985:IMP

[148] Mari Palta. Investigating maximum power losses in survival studies with nonstratified randomization. Biometrics, 41(2):497–504, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laake:1985:PMD

[149] Petter Laake, Knut Laake, and Rolf Aaberge. On the problem of measuring the distance between distribution functions: analysis of hospitalization versus mortality. Biometrics, 41(2):515–523, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cuzick:1985:MAC

[150] Jack Cuzick. A method for analysing case-control studies with ordinal exposure variables. Biometrics, 41(3):609–621, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mickey:1985:GMH

[151] Ruth M. Mickey and Robert M. Elashoff. A generalization of the Mantel– Haenszel estimator of partial association for 2×J ×K tables. Biometrics, 41(3):623–635, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Flanders:1985:NVE

[152] W. Dana Flanders. A new variance estimator for the Mantel–Haenszel odds ratio. Biometrics, 41(3):637–642, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [171].

Stablein:1985:TST

[153] Donald M. Stablein and I. A. Koutrouvelis. A two-sample test sensitive to crossing hazards in uncensored and singly censored data. Biometrics, 41 (3):643–652, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Soms:1985:SPI

[154] Andrew P. Soms. Simplified point and interval estimation for removal trapping. Biometrics, 41(3):663–668, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cowan:1985:BPR

[155] Richard Cowan. Branching process results in terms of moments of the generation-time distribution. Biometrics, 41(3):681–689, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Carter:1985:APL

[156] E. M. Carter and J. J. Hubert. Analysis of parallel-line assays with multivariate responses. Biometrics, 41(3):703–710, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Christensen:1985:GNM

[157] Erik R. Christensen and Chung-Yuan Chen. A general noninteractive multiple toxicity model including probit, logit, and Weibull transformations. Biometrics, 41(3):711–725, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonett:1985:SEL

[158] Douglas G. Bonett, J. Arthur Woodward, and P. M. Bentler. Some extensions of a linear model for categorical variables. Biometrics, 41(3): 745–750, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Golub:1985:SLS

[159] G. H. Golub. Solution of large sparse structured least-squares problem. Biometrics, 41(3):799, ???? 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OBrien:1985:CRT

[160] P. C. O’Brien and P. J. Dyck. Correction to: “A runs test based on run lengths” [Biometrics 41 (1985), no. 1, 237–244]. Biometrics, 41(3):821, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [145].

Sinclair:1985:TCD

[161] Dennis F. Sinclair, James E. Mosimann, and Duane A. Meeter. Tests for character displacement. Biometrics, 41(4):835–846, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sinclair:1985:TD

[162] Dennis F. Sinclair, James E. Mosimann, and Duane A. Meeter. Tests for deletion. Biometrics, 41(4):847–857, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Anderson:1985:GCM

[163] J. A. Anderson and J. D. Pemberton. The grouped continuous model for multivariate ordered categorical variables and covariate adjustment. Biometrics, 41(4):875–885, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pantula:1985:NAV

[164] Sastry G. Pantula and Kenneth H. Pollock. Nested analysis of variance with autocorrelated errors. Biometrics, 41(4):909–920, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finkelstein:1985:SMR

[165] Dianne M. Finkelstein and Robert A. Wolfe. A semiparametric model for regression analysis of interval-censored failure time data. Biometrics, 41 (4):933–945, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Grieve:1985:BAT

[166] A. P. Grieve. A Bayesian analysis of the two-period crossover design for clinical trials. Biometrics, 41(4):979–990, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [186].

Westfall:1985:SSS

[167] Peter Westfall. Simultaneous small-sample multivariate Bernoulli confidence intervals. Biometrics, 41(4):1001–1013, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Racine-Poon:1985:BAN

[168] A. Racine-Poon. A Bayesian approach to nonlinear random effects models. Biometrics, 41(4):1015–1023, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Shanmugam:1985:IPD

[169] Ramalingam Shanmugam. An intervened Poisson distribution and its medical application. Biometrics, 41(4):1025–1029, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hall:1985:CMI

[170] Peter Hall. Counting methods for inference in binary mosaics. Biometrics, 41(4):1049–1052, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Flanders:1985:CNV

[171] W. Dana Flanders. Correction: “A new variance estimator for the Mantel– Haenszel odds ratio” [Biometrics 41 (1985), no. 3, 637–642]. Biometrics, 41(4):1103, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [152].

Muenz:1985:CMM

[172] Larry R. Muenz and Lawrence V. Rubinstein. Correction: “Markov models for covariate dependence of binary sequences” [Biometrics 41 (1985), no. 1, 91–101]. Biometrics, 41(4):1103, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [139].

Stablein:1985:ATS

[173] Donald M. Stablein and I. A. Koutrouvelis. Addendum: “A two-sample test sensitive to crossing hazards in uncensored and singly censored data”. Biometrics, 41(4):1104, 1985. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sutherland:1986:MIM

[174] R. A. Sutherland. A method for inferring the minimum number of genes controlling the reactions of cultivars of a host plant species to isolates of a pathogen under a gene-for-gene model. Biometrics, 42(1):15–24, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sobel:1986:SNP

[175] Marc J. Sobel, Jonathan Arnold, and Milton Sobel. Statistics of natural populations. III. Sequential sampling plans for the estimation of gene frequencies. Biometrics, 42(1):45–65, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Srivastava:1986:MBC

[176] M. S. Srivastava. Multivariate bioassay, combination of bioassays, and Fieller’s theorem. Biometrics, 42(1):131–141, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [266].

Oakes:1986:ALP

[177] David Oakes. An approximate likelihood procedure for censored data. Biometrics, 42(1):177–182, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Munoz:1986:CPS

[178] Alvaro Muñoz and Bernard Rosner. Correction to: “Power and sample size for a collection of 2 × 2 tables” [Biometrics 40 (1984), no. 4, 995– 1004; MR0786176 (86d:62090)]. Biometrics, 42(1):229, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [127].

Robins:1986:EMH

[179] James Robins, Norman Breslow, and Sander Greenland. Estimators of the Mantel–Haenszel variance consistent in both sparse data and large-strata limiting models. Biometrics, 42(2):311–323, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dewanji:1986:NMS

[180] Anup Dewanji and J. D. Kalbfleisch. Nonparametric methods for survival/sacrifice experiments. Biometrics, 42(2):325–341, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jewell:1986:BCU

[181] Nicholas P. Jewell. On the bias of commonly used measures of association for 2×2 tables. Biometrics, 42(2):351–358, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kraemer:1986:MAS

[182] Helena Chmura Kraemer. A measure of 2 × 2 association with stable variance and approximately normal small-sample distribution: planning cost-effective studies. Biometrics, 42(2):359–370, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donner:1986:CCI

[183] Allan Donner and George Wells. A comparison of confidence interval methods for the intraclass correlation coefficient. Biometrics, 42(2):401– 412, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See corrections [204, 248].

Meisner:1986:CMB

[184] Morris Meisner, H. B. Kushner, and Eugene M. Laska. Combining multivariate bioassays. Biometrics, 42(2):421–427, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giesbrecht:1986:ADI

[185] F. G. Giesbrecht. Analysis of data from incomplete block designs. Biometrics, 42(2):437–448, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [273].

Grieve:1986:CBA

[186] A. P. Grieve. Correction to: “A Bayesian analysis of the two-period crossover design for clinical trials” [Biometrics 41 (1985), no. 4, 979–990]. Biometrics, 42(2):459, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [166].

Daudin:1986:SVM

[187] J.-J. Daudin. Selection of variables of mixed-variable discriminant analysis. Biometrics, 42(3):473–481, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Whittemore:1986:SEU

[188] Alice S. Whittemore and Joseph B. Keller. Survival estimation using splines. Biometrics, 42(3):495–506, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donner:1986:LSR

[189] Allan Donner and Walter W. Hauck. The large-sample relative efficiency of the Mantel–Haenszel estimator in the fixed-strata case. Biometrics, 42 (3):537–545, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Brownie:1986:CPC

[190] Cavell Brownie, James E. Hines, and James D. Nichols. Constantparameter capture–recapture models. Biometrics, 42(3):561–574, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1986:IEM

[191] D. A. Williams. Interval estimation of the median lethal dose. Biometrics, 42(3):641–645, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [249].

Brillinger:1986:NVV

[192] David R. Brillinger. The natural variability of vital rates and associated statistics. Biometrics, 42(4):693–734, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). With discussion and a reply by the author.

Breslow:1986:NVV

[193] N. Breslow, C. L. Chiang, J. M. Hoem, P. Jagers, N. Keiding, L. Kish, K. G. Manton, and J. W. Tukey. The natural variability of vital-rates and associated statistics — discussion. Biometrics, 42(4):712–732, ???? 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Tukey:1986:CNV

[194] John W. Tukey. Comments on The natural variability of vital rates and associated statistics. Biometrics, 42(??):729–732, ???? 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jensen:1986:SMS

[195] E. B. Jensen and R. Sundberg. Statistical models for stereological inference about spatial structures: on the applicability of best linear unbiased estimators in stereology. Biometrics, 42(4):735–751, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clayton:1986:SRD

[196] Gareth Clayton and Trevor F. Cox. Some robust density estimators for spatial point processes. Biometrics, 42(4):753–767, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jennrich:1986:URM

[197] Robert I. Jennrich and Mark D. Schluchter. Unbalanced repeatedmeasures models with structured covariance matrices. Biometrics, 42(4): 805–820, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fresen:1986:NFM

[198] John L. Fresen and June M. Juritz. A note on Foss’s method of obtaining initial estimates for exponential curve fitting by numerical integration.

Biometrics, 42(4):821–827, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finkelstein:1986:PHM

[199] Dianne M. Finkelstein. A proportional hazards model for interval-censored failure time data. Biometrics, 42(4):845–854, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buonaccorsi:1986:DSR

[200] John P. Buonaccorsi. Designs for slope ratio assays. Biometrics, 42(4): 875–882, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Simonoff:1986:AEP

[201] Jeffrey S. Simonoff, Yosef Hochberg, and Benjamin Reiser. Alternative estimation procedures for Pr(X < Y) in categorized data. Biometrics, 42 (4):895–907, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [317].

Loukas:1986:IDT

[202] S. Loukas and C. D. Kemp. The index of dispersion test for the bivariate Poisson distribution. Biometrics, 42(4):941–948, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Aebischer:1986:EPU

[203] N. J. Aebischer. Estimating the proportion of uncatchable animals in a population by double-sampling. Biometrics, 42(4):973–979, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donner:1986:CCC

[204] A. Donner and G. Wells. Correction to: “A comparison of confidence intervals for the intraclass correlation coefficient” [Biometrics 42 (1986), no. 2, 401–412]. Biometrics, 42(4):1009, 1986. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [183].

Helland:1987:IUR

[205] Inge S. Helland. On the interpretation and use of R^{2} in regression analysis. Biometrics, 43(1):61–69, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Halperin:1987:DFC

[206] Max Halperin, Peter R. Gilbert, and John M. Lachin. Distribution-free confidence intervals for Pr(X_{1} < X_{2}). Biometrics, 43(1):71–80, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

VanRyzin:1987:DRM

[207] John Van Ryzin and Kamta Rai. A dose-response model incorporating nonlinear kinetics. Biometrics, 43(1):95–105, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Mihalko:1987:RST

[208] Daniel Mihalko. Repeated significance testing in survival analysis using a litter-matched design. Biometrics, 43(1):115–125, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kalish:1987:EED

[209] Leslie A. Kalish and Colin B. Begg. Evaluation of efficient designs for observational epidemiologic studies. Biometrics, 43(1):145–167, 1987. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wang:1987:AOO

[210] Samuel K. Wang and Anastasios A. Tsiatis. Approximately optimal oneparameter boundaries for group sequential trials. Biometrics, 43(1):193– 199, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kemp:1987:RED

[211] A. W. Kemp. Review: Exploring Data Tables, Trends, and Shapes, by D. C. Hoaglin, F. Mosteller, and J. W. Tukey. Biometrics, 43(1):252, March 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

URL http://www.jstor.org/stable/2531968.

Barry:1987:ADB

[212] Daniel Barry and J. A. Hartigan. Asynchronous distance between homologous DNA sequences. Biometrics, 43(2):261–276, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Liang:1987:EMH

[213] Kung-Yee Liang. Extended Mantel–Haenszel estimating procedure for multivariate logistic regression models. Biometrics, 43(2):289–299, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Meng:1987:BAM

[214] Cliff Y. K. Meng and Arthur P. Dempster. A Bayesian approach to the multiplicity problem for significance testing with binomial data. Biometrics, 43(2):301–311, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lindsey:1987:MCA

[215] Jane C. Lindsey, Agnes M. Herzberg, and Donald G. Watts. A method for cluster analysis based on projections and quantile-quantile plots. Biometrics, 43(2):327–341, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1987:DFO

[216] Steven K. Thompson and Fred L. Ramsey. Detectability functions in observing spatial point processes. Biometrics, 43(2):355–362, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Taylor:1987:KSC

[217] Jeremy M. G. Taylor. Kendall’s and Spearman’s correlation coefficients in the presence of a blocking variable. Biometrics, 43(2):409–416, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holland:1987:ISR

[218] Burt S. Holland and Margaret DiPonzio Copenhaver. An improved sequentially rejective Bonferroni test procedure. Biometrics, 43(2):417–423, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [232].

Phillips:1987:EVM

[219] Abraham Phillips and Paul W. Holland. Estimators of the variance of the Mantel–Haenszel log-odds-ratio estimate. Biometrics, 43(2):425–431, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jannarone:1987:EIA

[220] Robert J. Jannarone, Caroline A. Macera, and Carol Z. Garrison. Evaluating interrater agreement through “case-control” sampling. Biometrics, 43 (2):433–437, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berry:1987:LTA

[221] Donald A. Berry. Logarithmic transformations in ANOVA. Biometrics, 43 (2):439–456, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Fuchs:1987:SCI

[222] Camil Fuchs and Allan R. Sampson. Simultaneous confidence intervals for the general linear model. Biometrics, 43(2):457–469, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Graubard:1987:CCS

[223] Barry I. Graubard and Edward L. Korn. Choice of column scores for testing independence in ordered 2 × K contingency tables. Biometrics, 43 (2):471–476, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pocock:1987:AME

[224] Stuart J. Pocock, Nancy L. Geller, and Anastasios A. Tsiatis. The analysis of multiple endpoints in clinical trials. Biometrics, 43(3):487–498, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lu:1987:CIR

[225] Tai-Fang C. Lu, Franklin A. Graybill, and Richard K. Burdick. Confidence intervals on the ratio of expected mean squares (θ_{1} + dθ_{2})/θ_{3}. Biometrics, 43(3):535–543, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Khuri:1987:ETM

[226] A. I. Khuri and R. C. Littell. Exact tests for the main effects variance components in an unbalanced random two-way model. Biometrics, 43(3): 545–560, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krzanowski:1987:CVP

[227] W. J. Krzanowski. Cross-validation in principal component analysis. Biometrics, 43(3):575–584, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thompson:1987:PSL

[228] E. A. Thompson and T. R. Meagher. Parental and sib likelihoods in genealogy reconstruction. Biometrics, 43(3):585–600, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:1987:MLE

[229] Philip J. Smith. Maximum likelihood estimation of the distribution of length at age. Biometrics, 43(3):601–615, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Vonesh:1987:EIR

[230] Edward F. Vonesh and Randy L. Carter. Efficient inference for randomcoefficient growth curve models with unbalanced data. Biometrics, 43(3): 617–628, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Johnson:1987:TSR

[231] Richard A. Johnson, Steve Verrill, and Dan H. Moore II. Two-sample rank tests for detecting changes that occur in a small proportion of the treated population. Biometrics, 43(3):641–655, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holland:1987:CIS

[232] B. S. Holland and M. D. Copenhaver. Correction to: “An improved sequentially rejective Bonferroni test procedure”. Biometrics, 43(3):737, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [218].

Brookmeyer:1987:BPC

[233] Ron Brookmeyer and Mitchell H. Gail. Biases in prevalent cohorts. Biometrics, 43(4):739–749, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wickramaratne:1987:CES

[234] Priya J. Wickramaratne and Theodore R. Holford. Confounding in epidemiologic studies: the adequacy of the control group as a measure of confounding. Biometrics, 43(4):751–765, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [331].

Chao:1987:EPS

[235] Anne Chao. Estimating the population size for capture–recapture data with unequal catchability. Biometrics, 43(4):783–791, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Raaijmakers:1987:SAM

[236] Jeroen G. W. Raaijmakers. Statistical analysis of the Michaelis–Menten equation. Biometrics, 43(4):793–803, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kunert:1987:VEC

[237] Joachim Kunert. On variance estimation in crossover designs. Biometrics, 43(4):833–845, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kim:1987:CIF

[238] Kyungmann Kim and David L. DeMets. Confidence intervals following group sequential tests in clinical trials. Biometrics, 43(4):857–864, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gross:1987:SSD

[239] Alan J. Gross, Hurshell H. Hunt, Alan B. Cantor, and B. Christine Clark. Sample size determination in clinical trials with an emphasis on exponentially distributed responses. Biometrics, 43(4):875–883, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Matts:1987:REZ

[240] John Matts and Richard McHugh. Randomization and efficiency in Zelen’s single-consent design. Biometrics, 43(4):885–894, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Crager:1987:ACP

[241] Michael R. Crager. Analysis of covariance in parallel-group clinical trials with pretreatment baselines. Biometrics, 43(4):895–901, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morgan:1987:NED

[242] Timothy M. Morgan. Nonparametric estimation of duration of accrual and total study length for clinical trials. Biometrics, 43(4):903–912, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Edwards:1987:ESB

[243] Don Edwards and Jack J. Berry. The efficiency of simulation-based multiple comparisons. Biometrics, 43(4):913–928, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Beal:1987:ACI

[244] S. L. Beal. Asymptotic confidence intervals for the difference between two binomial parameters for use with small samples. Biometrics, 43(4): 941–950, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bonett:1987:ESC

[245] Douglas G. Bonett, J. Arthur Woodward, and P. M. Bentler. Estimating the size of closed populations using inverse multiple-recapture sampling. Biometrics, 43(4):985–991, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Bedrick:1987:FCI

[246] Edward J. Bedrick. A family of confidence intervals for the ratio of two binomial proportions. Biometrics, 43(4):993–998, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Streit:1987:TEI

[247] Franz Streit. Testing for effective intervention. Biometrics, 43(4):999– 1000, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Donner:1987:CCC

[248] A. Donner and G. Wells. Correction to: “A comparison of confidence interval methods for the intraclass correlation coefficient” [Biometrics 42 (1986), no. 2, 401–412]. Biometrics, 43(4):1035, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [183].

Williams:1987:CIE

[249] D. A. Williams. Correction to: “Interval estimation of the median lethal dose” [Biometrics 42 (1986), no. 3, 641–645]. Biometrics, 43(4):1035, 1987. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [191].

Rice:1988:NPM

[250] William R. Rice. A new probability model for determining exact P -values for 2 × 2 contingency tables when comparing binomial proportions. Biometrics, 44(1):1–22, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krzanowski:1988:CDN

[251] W. J. Krzanowski and Y. T. Lai. A criterion for determining the number of groups in a data set using sum-of-squares clustering. Biometrics, 44 (1):23–34, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lee:1988:API

[252] Andy H. Lee. Assessing partial influence in generalized linear models. Biometrics, 44(1):71–77, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Self:1988:PSS

[253] Steven G. Self and Robert H. Mauritsen. Power/sample size calculations for generalized linear models. Biometrics, 44(1):79–86, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Buonaccorsi:1988:BIR

[254] John P. Buonaccorsi and Constantine A. Gatsonis. Bayesian inference for ratios of coefficients in a linear model. Biometrics, 44(1):87–101, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berger:1988:TCS

[255] Roger L. Berger, Dennis D. Boos, and Frank M. Guess. Tests and confidence sets for comparing two mean residual life functions. Biometrics, 44 (1):103–115, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lyons:1988:DPN

[256] N. I. Lyons and K. Hutcheson. Distributional properties of the number of moves index of diversity. Biometrics, 44(1):131–140, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Srivastava:1988:EII

[257] M. S. Srivastava, K. J. Keen, and R. S. Katapa. Estimation of interclass and intraclass correlations in multivariate familial data. Biometrics, 44 (1):141–150, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1988:ECC

[258] Margaret C. Wu and Raymond J. Carroll. Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process. Biometrics, 44(1):175–188, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [342].

Conover:1988:LMP

[259] William J. Conover and David S. Salsburg. Locally most powerful tests for detecting treatment effects when only a subset of patients can be expected to “respond” to treatment. Biometrics, 44(1):189–196, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:1988:MPL

[260] Peter F. Thall. Mixed Poisson likelihood regression models for longitudinal interval count data. Biometrics, 44(1):197–209, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [330].

Berry:1988:OSS

[261] Donald A. Berry and Chih-Hsiang Ho. One-sided sequential stopping boundaries for clinical trials: a decision-theoretic approach. Biometrics, 44(1):219–227, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lakatos:1988:SSB

[262] Edward Lakatos. Sample sizes based on the log-rank statistic in complex clinical trials. Biometrics, 44(1):229–241, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [281].

Thomas:1988:ACV

[263] D. Roland Thomas and J. N. K. Rao. On the analysis of categorical variables by linear models having singular covariance matrices. Biometrics, 44 (1):243–248, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Slud:1988:HDC

[264] E. Slud and D. Byar. How dependent causes of death can make risk factors appear protective. Biometrics, 44(1):265–269, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hall:1988:CIS

[265] Peter Hall. On confidence intervals for spatial parameters estimated from nonreplicated data. Biometrics, 44(1):271–277, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Srivastava:1988:CMB

[266] M. S. Srivastava. Correction to: “Multivariate bioassay, combination of bioassays, and Fieller’s theorem” [Biometrics 42 (1986), no. 1, 131–141; MR0845630 (87g:62161)]. Biometrics, 44(1):321, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [176].

Gart:1988:AIE

[267] John J. Gart and Jun mo Nam. Approximate interval estimation of the ratio of binomial parameters: a review and corrections for skewness. Biometrics, 44(2):323–338, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Kuritz:1988:ARE

[268] Stephen J. Kuritz and J. Richard Landis. Attributable risk estimation from matched case-control data. Biometrics, 44(2):355–367, 1988. CO-DEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Dinse:1988:ETI

[269] Gregg E. Dinse. Estimating tumor incidence rates in animal carcinogenicity experiments. Biometrics, 44(2):405–415, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heltshe:1988:JEM

[270] James F. Heltshe. Jackknife estimate of the matching coefficient of similarity. Biometrics, 44(2):447–460, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laska:1988:ESP

[271] Eugene M. Laska, Morris Meisner, and Carole Siegel. Estimating the size of a population from a single sample. Biometrics, 44(2):461–472, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Klein:1988:BNS

[272] John P. Klein and M. L. Moeschberger. Bounds on net survival probabilities for dependent competing risks. Biometrics, 44(2):529–538, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Giesbrecht:1988:CAD

[273] F. G. Giesbrecht. Correction to: “Analysis of data from incomplete block designs” [Biometrics 42 (1986), no. 2, 437–448]. Biometrics, 44(2):633, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [185].

Kitsos:1988:ODP

[274] C. P. Kitsos, D. M. Titterington, and B. Torsney. An optimal design problem in rhythmometry. Biometrics, 44(3):657–671, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Ball:1988:UAV

[275] Frank Ball and Peter Donnelly. A unified approach to variability in compartmental models. Biometrics, 44(3):685–694, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Cameron:1988:TCE

[276] M. A. Cameron, G. K. Eagleson, M. E. Willcox, D. G. Laing, and H. Panhuber. Testing for correlated errors in the analysis of variance. Biometrics, 44(3):695–704, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Coffey:1988:REM

[277] Mary Coffey. A random effects model for binary data from dependent samples. Biometrics, 44(3):787–801, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hirji:1988:EIM

[278] Karim F. Hirji, Cyrus R. Mehta, and Nitin R. Patel. Exact inference for matched case-control studies. Biometrics, 44(3):803–814, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Sylvester:1988:BAD

[279] Richard J. Sylvester. A Bayesian approach to the design of Phase II clinical trials. Biometrics, 44(3):823–836, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Williams:1988:ECI

[280] D. A. Williams. An exact confidence interval for the relative potency estimated from a multivariate bioassay. Biometrics, 44(3):861–867, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lakatos:1988:CSS

[281] E. Lakatos. Correction to: “Sample sizes based on the log-rank statistic in complex clinical trials” [Biometrics 44 (1988), no. 1, 229–241]. Biometrics, 44(3):923, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [262].

Diggle:1988:AAR

[282] Peter J. Diggle. An approach to the analysis of repeated measurements. Biometrics, 44(4):959–971, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Davis:1988:NMA

[283] Charles S. Davis and L. J. Wei. Nonparametric methods for analyzing incomplete nondecreasing repeated measurements. Biometrics, 44(4):1005– 1018, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeger:1988:MRM

[284] Scott L. Zeger and Bahjat Qaqish. Markov regression models for time series: a quasi-likelihood approach. Biometrics, 44(4):1019–1031, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Prentice:1988:CBR

[285] Ross L. Prentice. Correlated binary regression with covariates specific to each binary observation. Biometrics, 44(4):1033–1048, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeger:1988:MLD

[286] Scott L. Zeger, Kung-Yee Liang, and Paul S. Albert. Models for longitudinal data: a generalized estimating equation approach. Biometrics, 44(4): 1049–1060, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [303].

Knuiman:1988:IPI

[287] M. W. Knuiman and T. P. Speed. Incorporating prior information into the analysis of contingency tables. Biometrics, 44(4):1061–1071, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Morris:1988:SSC

[288] Max D. Morris. Small-sample confidence limits for parameters under inequality constraints with application to quantal bioassay. Biometrics, 44 (4):1083–1092, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Staniswalis:1988:KED

[289] Joan G. Staniswalis and Vanessa Cooper. Kernel estimates of dose response. Biometrics, 44(4):1103–1119, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thode:1988:SPP

[290] Henry C. Thode, Jr., Stephen J. Finch, and Nancy R. Mendell. Simulated percentage points for the null distribution of the likelihood ratio test for a mixture of two normals. Biometrics, 44(4):1195–1201, 1988. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

DeGruttola:1989:ADC

[291] Victor De Gruttola and Stephen W. Lagakos. Analysis of doubly-censored survival data, with application to AIDS. Biometrics, 45(1):1–11, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Gentleman:1989:DFO

[292] Jane F. Gentleman and Ronald C. Mullin. The distribution of the frequency of occurrence of nucleotide subsequences, based on their overlap capability. Biometrics, 45(1):35–52, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Hernandez:1989:DCA

[293] J. L. Hernández and B. S. Weir. A disequilibrium coefficient approach to Hardy–Weinberg testing. Biometrics, 45(1):53–70, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Clifford:1989:ASC

[294] Peter Clifford, Sylvia Richardson, and Denis Hémon. Assessing the significance of the correlation between two spatial processes. Biometrics, 45 (1):123–134, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

OQuigley:1989:STH

[295] John O’Quigley and Fabienne Pessione. Score tests for homogeneity of regression effect in the proportional hazards model. Biometrics, 45(1): 135–144, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Huster:1989:MPS

[296] William J. Huster, Ron Brookmeyer, and Steven G. Self. Modelling paired survival data with covariates. Biometrics, 45(1):145–156, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Jones:1989:GCN

[297] Michael P. Jones and John Crowley. A general class of nonparametric tests for survival analysis. Biometrics, 45(1):157–170, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Greenland:1989:GMH

[298] Sander Greenland. Generalized Mantel–Haenszel estimators for K 2 × J tables. Biometrics, 45(1):183–191, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rochon:1989:MLE

[299] James Rochon and Ronald W. Helms. Maximum likelihood estimation for incomplete repeated measures experiments under an ARMA covariance structure. Biometrics, 45(1):207–218, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Finkelstein:1989:AMT

[300] Dianne M. Finkelstein and David A. Schoenfeld. Analysis of multiple tumor data from a rodent carcinogenicity experiment. Biometrics, 45(1): 219–230, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chang:1989:CIN

[301] Myron N. Chang. Confidence intervals for a normal mean following a group sequential test. Biometrics, 45(1):247–254, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Snapinn:1989:EER

[302] Steven M. Snapinn and James D. Knoke. Estimation of error rates in discriminant analysis with selection of variables. Biometrics, 45(1):289–299, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zeger:1989:CML

[303] S. L. Zeger, K.-Y. Liang, and P. S. Albert. Correction to: “Models for longitudinal data: a generalized estimating equation approach” [Biometrics 44 (1988), no. 4, 1049–1060]. Biometrics, 45(1):347, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [286].

Rudemo:1989:REM

[304] M. Rudemo, D. Ruppert, and J. C. Streibig. Random-effect models in nonlinear regression with applications to bioassay. Biometrics, 45(2):349–362, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Roeder:1989:AML

[305] Kathryn Roeder, B. Devlin, and Bruce G. Lindsay. Application of maximum likelihood methods to population genetic data for the estimation of individual fertilities. Biometrics, 45(2):363–379, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Moulton:1989:ARM

[306] Lawrence H. Moulton and Scott L. Zeger. Analyzing repeated measures on generalized linear models via the bootstrap. Biometrics, 45(2):381–394, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Steinhorst:1989:SAM

[307] R. Kirk Steinhorst and Michael D. Samuel. Sightability adjustment methods for aerial surveys of wildlife populations. Biometrics, 45(2):415–425, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Chao:1989:EPS

[308] Anne Chao. Estimating population size for sparse data in capture– recapture experiments. Biometrics, 45(2):427–438, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Heimbuch:1989:CRE

[309] Douglas G. Heimbuch and John M. Hoenig. Change-in-ratio estimators for habitat usage and relative population size. Biometrics, 45(2):439–451, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Zahl:1989:LTS

[310] Samuel Zahl. Line transect sampling with unknown probability of detection along the transect. Biometrics, 45(2):453–470, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Yip:1989:IPC

[311] Paul Yip. An inference procedure for a capture and recapture experiment with time-dependent capture probabilities. Biometrics, 45(2):471–479, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laird:1989:EBC

[312] Nan M. Laird and Thomas A. Louis. Empirical Bayes confidence intervals for a series of related experiments. Biometrics, 45(2):481–495, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pepe:1989:WKM

[313] Margaret Sullivan Pepe and Thomas R. Fleming. Weighted Kaplan–Meier statistics: a class of distance tests for censored survival data. Biometrics, 45(2):497–507, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Halperin:1989:DFC

[314] Max Halperin, Mohamed I. Hamdy, and Peter F. Thall. Distribution-free confidence intervals for a parameter of Wilcoxon–Mann–Whitney type for ordered categories and progressive censoring. Biometrics, 45(2):509–521, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:1989:TSD

[315] Peter F. Thall, Richard Simon, and Susan S. Ellenberg. A two-stage design for choosing among several experimental treatments and a control in clinical trials. Biometrics, 45(2):537–547, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Rosenbaum:1989:RKE

[316] Paul R. Rosenbaum. The role of known effects in observational studies. Biometrics, 45(2):557–569, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Simonoff:1989:CAE

[317] J. S. Simonoff, Y. Hochberg, and B. Reiser. Correction to: “Alternative estimation procedures for Pr(X < Y) in categorized data” [Biometrics 42 (1986), no. 4, 895–907; MR0872965 (88c:62045)]. Biometrics, 45(2):701, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [201].

Matis:1989:SGS

[318] James H. Matis, Thomas E. Wehrly, and William C. Ellis. Some generalized stochastic compartment models for digesta flow. Biometrics, 45(3): 703–720, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Lwin:1989:PM

[319] T. Lwin and P. J. Martin. Probits of mixtures. Biometrics, 45(3):721–732, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laurence:1989:OSM

[320] A. F. Laurence and B. J. T. Morgan. Observations on a stochastic model for quantal assay data. Biometrics, 45(3):733–744, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Izenman:1989:CLS

[321] Alan J. Izenman and J. S. Williams. A class of linear spectral models and analyses for the study of longitudinal data. Biometrics, 45(3):831–849, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Raz:1989:ARM

[322] Jonathan Raz. Analysis of repeated measurements using nonparametric smoothers and randomization tests. Biometrics, 45(3):851–871, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Smith:1989:ALS

[323] Woollcott Smith. ANOVA-like similarity analysis using expected species shared. Biometrics, 45(3):873–881, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Emerson:1989:SGS

[324] Scott S. Emerson and Thomas R. Fleming. Symmetric group sequential test designs. Biometrics, 45(3):905–923, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Storer:1989:DAP

[325] Barry E. Storer. Design and analysis of Phase I clinical trials. Biometrics, 45(3):925–937, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1989:ECC

[326] Margaret C. Wu and Kent R. Bailey. Estimation and comparison of changes in the presence of informative right censoring: conditional linear model. Biometrics, 45(3):939–955, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See correction [341].

Beal:1989:SSD

[327] S. L. Beal. Sample size determination for confidence intervals on the population mean and on the difference between two population means. Biometrics, 45(3):969–977, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Choi:1989:RFG

[328] Jai Won Choi and Richard B. McHugh. A reduction factor in goodness-of-fit and independence tests for clustered and weighted observations. Biometrics, 45(3):979–996, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Holland:1989:NCM

[329] Paul W. Holland. A note on the covariance of the Mantel–Haenszel logodds-ratio estimator and the sample marginal rates. Biometrics, 45(3): 1009–1016, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Thall:1989:CMP

[330] Peter F. Thall. Correction to: “Mixed Poisson likelihood regression models for longitudinal interval count data” [Biometrics 44 (1988), no. 1, 197– 209]. Biometrics, 45(3):1039, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [260].

Wickramaratne:1989:CCE

[331] Priya J. Wickramaratne and Theodore R. Holford. Correction to: “Confounding in epidemiologic studies: the adequacy of the control group as a measure of confounding” [Biometrics 43 (1987), no. 4, 751–765]. Biometrics, 45(3):1039, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [234].

Robins:1989:DSC

[332] James M. Robins, Ross L. Prentice, and Donald Blevins. Designs for synthetic case-control studies in open cohorts. Biometrics, 45(4):1103–1116, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Schell:1989:IFR

[333] Michael J. Schell and Frederick W. Leysieffer. An increasing failure rate approach to low-dose extrapolation. Biometrics, 45(4):1117–1123, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Robins:1989:PCU

[334] James Robins and Sander Greenland. The probability of causation under a stochastic model for individual risk. Biometrics, 45(4):1125–1138, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Laska:1989:TWI

[335] Eugene M. Laska and Morris J. Meisner. Testing whether an identified treatment is best. Biometrics, 45(4):1139–1151, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Pettitt:1989:ODF

[336] A. N. Pettitt. One degree of freedom for nonadditivity: applications with generalized linear models and link functions. Biometrics, 45(4):1153–1162, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Krzanowski:1989:NCT

[337] W. J. Krzanowski and D. Radley. Nonparametric confidence and tolerance regions in canonical variate analysis. Biometrics, 45(4):1163–1173, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Berry:1989:MAD

[338] Donald A. Berry. Monitoring accumulating data in a clinical trial. Biometrics, 45(4):1197–1211, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Marshall:1989:PVS

[339] R. J. Marshall. The predictive value of simple rules for combining two diagnostic tests. Biometrics, 45(4):1213–1222, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

McDonald:1989:IES

[340] J. F. McDonald and D. Palanacki. Interval estimation of the size of a small population from a mark-recapture experiment. Biometrics, 45(4): 1223–1231, 1989. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic).

Wu:1990:CEC

[341] M. C. Wu and K. R. Bailey. Correction to: “Estimation and comparison of changes in the presence of informative right censoring: conditional linear model” [Biometrics 45 (1989), no. 3, 939–955; MR1029611 (90k:62151)]. Biometrics, 46(3):889, 1990. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [326].

Wu:1991:CEC

[342] M. C. Wu and R. J. Carroll. Correction to: “Estimation and comparison of changes in the presence of informative right censoring by modeling the censoring process” [Biometrics 44 (1988), no. 1, 175–188; MR0931633 (89d:62109)]. Biometrics, 47(1):357, 1991. CODEN BIOMB6. ISSN 0006-341X (print), 1541-0420 (electronic). See [258].